# GradNetOT: Learning Optimal Transport Maps with GradNets

### v1

@@ VI-B Robot Swarms as Optimal Transport
The first column corresponds to the transport map learned by the mGradNet-M. The second column corresponds to the transport map from the barycentric projection. The barycentric projection converts a discrete optimal transport plan, from solving the Monge-Kantarovich relaxation, to an optimal transport map. Given an optimal transport plan $\gamma\in\mathbb{R}^{a\times b}$, which describes how much mass to move from source points $x_{i}$ to target points $y_{j}$, the barycentric projection described in [13] is defined as:
$T(x_{i})=\mathbb{E}_{\gamma}[Y|X=x_{i}]=\frac{\sum_{j=1}^{b}\gamma_{ij}y_{j}}{\sum_{j=1}^{b}\gamma_{ij}}$ (10)
Specifically, the barycentric projection approximates a Monge (not mass splitting) transport map using a potentially mass-splitting Kantorovich plan. When the solution to the Kantarovich problem $\gamma$ does not split mass, and a Monge map exists, then the barycentric projection recovers the optimal Monge map in the limit as the number of source and target points increases [13]. In Table I we report the mean squared error (MSE) between the barycentric projection $T(x_{i})$ and the learned mapping $T_{\theta}(x_{i})$ over 1000 samples $x_{i}$ for various source and target image pairs.
@@
Each entry in the table reports the difference between the transport map learned by an mGradNet-M and the barycentric projection, when transporting mass from the source density (row) to the target density (column). Overall, we observe both numerically and qualitatively in Fig. 3, that our proposed method closely matches the performance of the barycentric projection.
@@ VII Conclusion
We propose a method for learning optimal transport maps using mGradNets – neural networks guaranteed to correspond to gradients of convex functions. Inspired by Brenier’s theorem, our approach finds solutions to the Monge problem with quadratic Euclidean cost by directly searching over the space of gradients of convex potentials by optimizing a Monge-Ampère loss. We validate our method through a series of experiments demonstrating its effectiveness. Our framework opens avenues for further research in the design and training of theoretically-motivated neural networks for learning optimal transport maps.

### v2

@@ VI-B Image Morphing as Optimal Transport
The first column corresponds to the transport map learned by the mGradNet-M. The second column corresponds to the transport map from the barycentric projection. The barycentric projection converts a discrete optimal transport plan, from solving the Monge-Kantorovich relaxation, to an optimal transport map. Given an optimal transport plan $\gamma\in\mathbb{R}^{N_{x}\times N_{y}}$, which describes how much mass to move from source points $\{{\mathbf{x}}_{i}\}_{i=1}^{N_{x}}$ to target points $\{{\mathbf{y}}_{j}\}_{j=1}^{N_{y}}$, the barycentric projection described in [13] is defined as:
$T({\mathbf{x}}_{i})=\frac{\sum_{j=1}^{N_{y}}\gamma_{ij}{\mathbf{y}}_{j}}{\sum_{j=1}^{N_{y}}\gamma_{ij}}$ (11)
Specifically, the barycentric projection approximates a Monge (not mass splitting) transport map using a potentially mass-splitting Kantorovich plan. When the solution to the Kantorovich problem $\gamma$ does not split mass, and a Monge map exists, then the barycentric projection recovers the optimal Monge map in the limit as the number of source and target points increases [13, 25]. However, computing the optimal transport plan scales quadratically with the number of points, motivating the learning of a parametric OT map, which avoids the cost of solving and storing a full discrete transport plan.
@@
In Table I we report the mean squared error (MSE) between the barycentric projection $T({\mathbf{x}}_{i})$ and the learned mapping $T_{\boldsymbol{\theta}}({\mathbf{x}}_{i})$ over 1000 samples ${\mathbf{x}}_{i}$ for various source and target image pairs. Each entry in the table reports the difference between the transport map learned by an mGradNet-M and the barycentric projection, when transporting mass from the source density (row) to the target density (column). Lower MSE values in the table imply that the transport cost of the learned OT map is closer to that of the optimal transport map. Overall, we observe both numerically and qualitatively in Fig. 3, that our proposed method closely matches the performance of the barycentric projection.
@@ VII Conclusion
We propose a method for learning optimal transport maps using mGradNets – neural networks guaranteed to correspond to gradients of convex functions. Inspired by Brenier’s theorem, our approach finds solutions to the Monge problem with quadratic Euclidean cost by directly searching over the space of gradients of convex potentials by optimizing a Monge-Ampère loss. We validate our method through a series of experiments demonstrating its effectiveness. Our framework opens avenues for further research in the design and training of theoretically-motivated neural networks for solving optimal transport problems.